# Dissecting users’ needs for search result explanations

### 29 Jan 2024 v1

@@ 1. Introduction
Search engines have become an integral part of our daily lives. According to a 2023 report, an overwhelming 93% of all web traffic flows through search engines (M., 2023). They not only facilitate information dissemination but also exert a profound influence on our beliefs, shape ideas, and mold the behaviors and perceptions of users (Carroll, 2014). For instance, a past study revealed that partisan bias in search engine rankings can substantially influence the voting intentions of undecided voters (Epstein et al., 2017; Epstein and Robertson, 2015). Despite the power they hold, search engines have long operated as black boxes, providing little to no explanation about why specific search results are displayed to users and the order in which they appear. This lack of transparency has led users to develop folk theories or incorrect concepts about search engine functioning. For instance, a previous study revealed that searchers mistakenly believe web search engines accept payment for higher-ranking positions in search results, beyond advertisements (Thomas et al., 2019). The lack of transparency has also been linked to limited engagement with search results (Muramatsu and Pratt, 2001) and has been shown to contribute to users’ difficulty in finding desired information (Thomas et al., 2019; Holman, 2011; Eslami et al., 2016). Therefore, researchers and scholars have advocated for making search algorithms more accountable by being more transparent and interpretable. Studies have shown that by explaining how search works and communicating more information about search results, we can help users be more efficient and effective searchers (Ramos and Eickhoff, 2020; Koenemann and Belkin, 1996; Muramatsu and Pratt, 2001; Hamilton et al., 2014) as well as raise user awareness of biases in search engine results (Munson et al., 2013).
@@
In this study, we also assess the utility of current search engine explanation features. Popular web search engines like Google and Bing have traditionally offered brief text snippets with each search result, highlighting relevant keywords to indicate content alignment with queries. However, in response to the growing demand for transparency and accountability, modern search engines have recently introduced more sophisticated features designed to provide search explanations and additional context for search results. there is a limited understanding of users’ awareness and the perceived effectiveness of these additional features. Our research contributes by investigating users’ perceptions of their usefulness. Overall, this study is guided by three fundamental research questions:
@@ 2.1. Users’ Interaction with Search Engines: Behavior, Decision-Making, and Mental Models
Scholars have also examined users’ mental models of online search. Studies in this area have sought to understand how users conceptualize search engines and the mechanisms behind them. Hendry and Efthimiadis, for instance, solicited sketches from students on how search engines work (Hendry and Efthimiadis, 2008). The resulting figures contained a diverse array of concepts along with several misconceptions about how search engines work (Hendry and Efthimiadis, 2008). Maramutsu et al conducted a user study to understand users’ knowledge of and reactions to the query transformations employed by search engines like stop word removals, boolean operators, etc. (Muramatsu and Pratt, 2001). Zhang et al. studied undergraduate students’ comprehension of various aspects of search engines, including search engine components, the search process, search result ranking mechanisms, and attributes of search engines (Zhang, 2008). Thomas et al combined surveys with interviews to understand how users conceptualize search engines’ ranking algorithms (Zhang, 2008). The authors aimed to identify the concepts about how a search engine operates that users are familiar with and the concepts that are alien suggesting that search engines can add the concepts to the explanations about search results about which they have little understanding (Thomas et al., 2019). In contrast, our approach takes a different trajectory. Instead of pinpointing specific concepts to add to explanations, we investigate the necessity of explanations themselves and the situations in which they prove beneficial.
@@ 2.2. Study context: explanations
The concept of explanation is multifaceted, drawing insights from various academic disciplines; however, there is no universally agreed-upon definition or formalization (Islam et al., 2020). According to psychologist Lombrozo, explanations serve as the “currency in which we exchange belief”. They encompass both a cognitive process, wherein causes for an event, including specific counterfactual cases, are identified, and a resulting product—an explanation—derived from the cognitive explanation process (Lombrozo, 2006). Numerous cognitive science theories emphasize that explanations frequently involve causal relations (Lewis, 1986; Srinivasan and Chander, 2021) and are supported by the idea that explaining an event entails offering information about its causal history. This notion aligns with the philosophical perspective that scientific explanations (Salmon, 1984; Woodward, 2005) predominantly rely on causal relationships, wherein scientific facts are elucidated by appealing to their causes. Explanations are also thought of as similar to proofs in logic where a set of fundamental laws is presented as axioms, and the deductive sequences within the proofs form the explanation. (Hempel and Oppenheim, 1948). However, these methods struggle to explain the mechanism of phenomena and may not be easily understandable for non-experts (Srinivasan and Chander, 2021).
Some definitions of explanation are specific to AI/ML and also take different forms. Some view it as any interpretable version of the original model (Lundberg and Lee, 2017). Others define it as the set of features in the interpretable domain influencing a decision for a specific example (Montavon et al., 2018). Additionally, a broader perspective encompasses all explanatory details about the model, including aspects like training data, performance, uncertainty, etc. (Liao et al., 2020). In the field of human-computer interaction, an explanation serves as a bridge between humans and a decision maker. It acts as a dual-purpose interface, accurately representing the decision maker while also being understandable to humans (Guidotti et al., 2018). Scholars highlight the social significance of explanations, going beyond the perspectives of developers and researchers. They define a ‘good explanation’ as one that is not only understood by the explainer but also holds meaning for the person seeking it (Brandão et al., 2019).
@@ 2.3. Search result explanations
There exists a substantial body of scholarly work dedicated to the design and generation of explanations (see (Abdul et al., 2018) for a detailed review). Previous research has focused on determining users’ needs from explanations (Liao et al., 2020; Lim and Dey, 2009; Lim et al., 2009), determining the impact of explanations on user perceptions of and resulting interactions with these systems (Kunkel et al., 2019; Cai et al., 2019; Cheng et al., 2019b; Dodge et al., 2019; Kocielnik et al., 2019; Kizilcec, 2016; Lim et al., 2009; Pu and Chen, 2006; Cheng et al., 2019b; Cramer et al., 2008), as well as developing metrics to evaluate the quality and effectiveness of explanations (Hoffman et al., 2018; Carvalho et al., 2019). However, most of this research primarily focuses on the domains of AI and ML, where the primary goal of explanations is to demystify how the model’s output is generated (Anik and Bunt, 2021; Letham et al., 2015; Ribeiro et al., 2016; Millecamp et al., 2019; Wang et al., 2023; Sendak et al., 2020) or within collaborative filtering and recommender systems, where explanations shed light on the rationale behind the generation of specific recommendations (Cheng et al., 2019a; Herlocker et al., 2000; Vig et al., 2009; Tintarev and Masthoff, 2010). In contrast, studying explanations for search engines is still an under-explored area in the broader field of explanations research. This gap in research is particularly noteworthy given the unparalleled impact of search engines, facilitating access to an extensive array of information spanning educational research to health inquiries, setting them apart significantly from ML or recommendation systems that often focus on specific content domains. The impact of search engines is evident with the results of recent surveys conducted by Edelman and Reuters that showed users trusting search engines more than any other source including traditional news outlets and news recommended on social media (Edelman, 2019; Newman et al., 2020). Given their impact and concerns regarding potential biases (Metaxa et al., 2021; Robertson et al., 2018; Hussein et al., 2020; Hu et al., 2019; Kay et al., 2015), scholars have begun advocating for transparent and explainable search interfaces.
@@
It is crucial to note that, while both search engines and recommendation systems operate as black boxes, their operational distinctions are profound. Search engines, propelled by explicit user queries, provide immediate and diverse results that hold the power to directly influence explicit decisions. In contrast, recommendation systems rely on implicit user behavior, gradually shaping decisions over time. Therefore transparency initiatives in the recommendation or AI/ML domains may fall short of adequately addressing the intricate interplay of user intent, query context, and the varied content types that search engines navigate. Thus, there is ongoing research focused solely on generating explanations for search results (Yu et al., 2022; Polley et al., 2021; Baier and Kaisers, 2020; Ai et al., 2019; Ai and Narayanan. R, 2021) and showing the impact of explanations on efficiency and trust (Ramos and Eickhoff, 2020; White et al., 2009; Eickhoff et al., 2014), as well as explaining the relevance of search results to users (Coyle and Smyth, 2005; Hearst, 1995). However, there is still a limited understanding of if and when users find search explanations beneficial. Our study explores the potential advantages of search explanations, identify the search objectives in which they prove valuable and ascertain the specific characteristics desired in valuable explanations.
@@ 3.1.2. Targeted user group
No. Search objective Example search queries Helpfulness of explanations
@@ 3.2.2. Study survey
Figure 1. (a) Figure showing highlighted keywords in the snippet of text accompanying the search result, offering insights into the alignment of the search result with the user’s search query, (b) Figure showing search explanations provided as an example during the interview. This example illustrates an overall explanation describing how the search engine works and, in addition, explains how relevance influenced the ranking of search results.
@@ 3.3.2. Interview protocol
Participants were first asked general warm-up questions, including the frequency of their search engine usage and whether they had encountered search explanations on the search results page. They were then prompted to recall situations where they questioned the search results returned by the search engines. Next, we conducted a contextual inquiry task, wherein participants were instructed to search for a specific query on their preferred search engine and share their screen. They were asked to describe their process for selecting search results and provide suggestions for additional information that could enhance their search experience. For this task, we selected two search queries voluntarily provided by participants for various search scenarios in the study survey. These included queries about both non-controversial topics such as ’best ways to learn investment,’, ’date ideas in Philly’, etc. and more controversial subjects (if such examples were provided by the participants) such as ’covid-19 vaccines,’ ’climate change’, etc. Recall that participants had initially selected these search queries from their browser histories during the study survey, mirroring real-world scenarios where they actively formulate search inquiries based on their information needs. Our goal was to ensure that the chosen queries resonated with participants’ familiarity and interests. During the interview process, some participants chose to search for additional or different search queries in addition to the ones we selected for them. We allowed this flexibility to ensure participants’ comfort and autonomy. Note that the objective of this task is not to determine specific search features needed in particular scenarios but rather to capture the range of potential features that could enhance participants’ overall search experience.
@@ 5.1. Utility of explanations in various search objectives
Table 3. Table providing a summary of the perceived benefits of search result explanations as expressed by the participants along with an example quote
@@ 5.2.2. Explanations could serve as decision-making tools by assisting in the selection of search results:
Table 4.
@@ 6. RQ3: Users’ perception of search explanations and search context provided by search engines
Option to access additional information is valued: All participants acknowledged the value of having the option to access additional information about the search result in both Bing and Google search engines. As P7 noted, “That gives a person the searcher the option of taking the extra time to learn the extra information… and that’s always a good thing to have”. Participants also appreciated the concise presentation of the information presented by search engines, as one participant mentioned, “I like that it’s concise and doesn’t show you everything you want to see..you have to click somewhere else, which I do like because it’s like it makes it less cluttered I guess” (P6).
@@
Broad and obvious explanations are not deemed useful: Many participants did not find the content of the search explanations, currently only provided by Google, particularly useful, deeming it to be either obvious or too broad. P5 provided an example stating that their getting a specific result due to their location in the USA is not very helpful—“I don’t think the result in the region would be particularly helpful, especially if it’s as broad as the United States”. P2 concurred, expressing, “I feel like a lot of this information doesn’t necessarily help me out with some of those qualities that we were talking about..like the trust, the credibility, the efficiency”.
@@
Contestability is desired in search: The majority of participants appreciated the functionality provided solely by Google to remove search results and offer feedback.. They saw this feature as a means to contest and influence the search results, granting them a voice in determining the relevance and accuracy of the displayed information. Participants appreciated the agency and empowerment that came with being able to shape the search results according to their preferences and needs. One participant emphasized the significance of user input, expressing the necessity of having such a feature—“I feel like it’s something that’s necessary to have something like this because if they did not include like a send feedback option, it would feel sort of wrong in a way, it would feel more so that like Google is saying, we know what’s right, we know what’s best, and you know, you don’t have a say in that” (P6).
@@
Privacy and personalization settings: Participants found the privacy and personalization settings provided only by Google to be helpful. The majority of them were not aware that they could change their privacy settings to alter the way Google collects their personal information to personalize the search results for them. Participants suggested that instead of having the privacy settings within the search explanations pop up, “it should be easily accessible and visualizable on the search result page itself” (P3).

### 23 Feb 2024 v2

@@ 1. Introduction
Search engines have become an integral part of our daily lives. According to a 2023 report, an overwhelming 93% of all web traffic flows through search engines (M., 2023). They not only facilitate information dissemination but also exert a profound influence on our beliefs, shape ideas, and mold the behaviors and perceptions of users (Carroll, 2014). For instance, a past study revealed that partisan bias in search engine rankings can substantially influence the voting preferences of individuals yet to make a decision (Epstein et al., 2017; Epstein and Robertson, 2015). Despite the power they hold, search engines have long operated as black boxes, providing little to no explanation about why specific search results are displayed to users and the order in which they appear. This lack of transparency has led users to develop folk theories or incorrect concepts about search engine functioning. For instance, a previous study revealed that searchers mistakenly believe web search engines accept payment for higher-ranking positions in search results, beyond advertisements (Thomas et al., 2019). The lack of transparency has also been linked to limited engagement with search results (Muramatsu and Pratt, 2001) and has been shown to contribute to users’ difficulty in finding desired information (Thomas et al., 2019; Holman, 2011; Eslami et al., 2016). Therefore, researchers and scholars have advocated for making search algorithms more accountable by being more transparent and interpretable. Studies have shown that by explaining how search works and communicating more information about search results, we can help users be more efficient and effective searchers (Ramos and Eickhoff, 2020; Koenemann and Belkin, 1996; Muramatsu and Pratt, 2001; Hamilton et al., 2014) as well as raise user awareness of biases in search engine results (Munson et al., 2013).
@@
In this study, we also assess the utility of current search engine explanation features. Popular web search engines like Google and Bing have traditionally offered brief text snippets with each search result, highlighting relevant keywords to indicate content alignment with queries. In response to the growing demand for transparency and accountability, modern search engines have recently introduced more sophisticated features designed to provide search explanations and additional context for search results. However, there is a limited understanding of users’ awareness and the perceived effectiveness of these additional features. Our research endeavors to bridge this gap by investigating users’ perceptions of their usefulness. Overall, this study is guided by three fundamental research questions:
@@ 2.1. Users’ Interaction with Search Engines: Behavior, Decision-Making, and Mental Models
Scholars have also examined users’ mental models of online search. Studies in this area have sought to understand how users conceptualize search engines and the mechanisms behind them. Hendry and Efthimiadis, for instance, solicited sketches from students on how search engines work (Hendry and Efthimiadis, 2008). The resulting figures contained a diverse array of concepts along with several misconceptions about how search engines work (Hendry and Efthimiadis, 2008). Maramutsu et al. conducted a user study to understand users’ knowledge of and reactions to the query transformations employed by search engines like stop word removals, boolean operators, etc. (Muramatsu and Pratt, 2001). Zhang et al. studied undergraduate students’ comprehension of various aspects of search engines, including search engine components, the search process, search result ranking mechanisms, and attributes of search engines (Zhang, 2008). Thomas et al. combined surveys with interviews to understand how users conceptualize search engines’ ranking algorithms (Zhang, 2008). The authors aimed to identify the concepts about how a search engine operates that users are familiar with and the concepts that are alien suggesting that search engines can add the concepts to the explanations about search results about which they have little understanding (Thomas et al., 2019). In contrast, our approach takes a different trajectory. Instead of pinpointing specific concepts to add to explanations, we investigate the necessity of explanations themselves and the situations in which they prove beneficial.
@@ 2.2. Study context: explanations
The concept of explanation is multifaceted, drawing insights from various academic disciplines; however, there is no universally agreed-upon definition or formalization (Islam et al., 2020). According to psychologist Lombrozo, explanations serve as the “currency in which we exchange belief”. They encompass both a cognitive process, wherein causes for an event, including specific counterfactual cases, are identified, and a resulting product—an explanation—derived from the cognitive explanation process (Lombrozo, 2006). Numerous cognitive science theories emphasize that explanations frequently involve causal relations (Lewis, 1986; Srinivasan and Chander, 2021) and are supported by the idea that explaining an event entails offering information about its causal history. This notion aligns with the philosophical perspective that scientific explanations (Salmon, 1984; Woodward, 2005) predominantly rely on causal relationships, wherein scientific facts are elucidated by appealing to their causes. Explanations are also thought of as similar to proofs in logic where a set of fundamental laws is presented as axioms, and the deductive sequences within the proofs form the explanation. (Hempel and Oppenheim, 1948). However, non-experts often struggle to understand these methods (Srinivasan and Chander, 2021).
Some definitions of explanation are specific to AI/ML and also take different forms. Some view it as any interpretable version of the original model (Lundberg and Lee, 2017). Others define it as the set of features in the interpretable domain influencing a decision for a specific example (Montavon et al., 2018). Additionally, a broader perspective encompasses all explanatory details about the model, including aspects like training data, performance, uncertainty, etc. (Liao et al., 2020). In the field of human-computer interaction, an explanation serves as a bridge between humans and a decision maker. It acts as a dual-purpose interface, accurately representing the decision maker while also being understandable to humans (Guidotti et al., 2018). Scholars also highlight the social significance of explanations, going beyond the perspectives of developers and researchers. They define a ‘good explanation’ as one that is not only understood by the explainer but also holds meaning for the person seeking it (Brandão et al., 2019).
@@ 2.3. Search result explanations
There exists a substantial body of scholarly work dedicated to the design and generation of explanations (see (Abdul et al., 2018) for a detailed review). Previous research has focused on determining users’ needs from explanations (Liao et al., 2020; Lim and Dey, 2009; Lim et al., 2009), determining the impact of explanations on user perceptions of and resulting interactions with these systems (Kunkel et al., 2019; Cai et al., 2019; Cheng et al., 2019b; Dodge et al., 2019; Kocielnik et al., 2019; Kizilcec, 2016; Lim et al., 2009; Pu and Chen, 2006; Cheng et al., 2019b; Cramer et al., 2008), as well as developing metrics to evaluate the quality and effectiveness of explanations (Hoffman et al., 2018; Carvalho et al., 2019). However, most of this research primarily focuses on the domains of AI and ML, where the primary goal of explanations is to demystify how the model’s output is generated (Anik and Bunt, 2021; Letham et al., 2015; Ribeiro et al., 2016; Millecamp et al., 2019; Wang et al., 2023; Sendak et al., 2020) or within collaborative filtering and recommender systems, where explanations shed light on the rationale behind the generation of specific recommendations (Cheng et al., 2019a; Herlocker et al., 2000; Vig et al., 2009; Tintarev and Masthoff, 2010). In contrast, studying explanations for search engines is still an under-explored area in the broader field of explanations research. This gap in research is particularly noteworthy given the unparalleled impact of search engines, facilitating access to an extensive array of information spanning educational research to health inquiries, setting them apart significantly from ML or recommendation systems that often focus on specific content domains. The impact of search engines is also evident with the results of recent surveys conducted by Edelman and Reuters that showed users trusting search engines more than any other source including traditional news outlets and news recommended on social media (Edelman, 2019; Newman et al., 2020). Given their impact and concerns regarding potential biases (Metaxa et al., 2021; Robertson et al., 2018; Hussein et al., 2020; Hu et al., 2019; Kay et al., 2015), scholars have begun advocating for transparent and explainable search interfaces.
@@
It is also crucial to note that, while both search engines and recommendation systems operate as black boxes, their operational distinctions are profound. Search engines, propelled by explicit user queries, provide immediate and diverse results that hold the power to directly influence explicit decisions. In contrast, recommendation systems rely on implicit user behavior, gradually shaping decisions over time. Therefore transparency initiatives in the recommendation or AI/ML domains may fall short of adequately addressing the intricate interplay of user intent, query context, and the varied content types that search engines navigate. Thus, there is ongoing research focused solely on generating explanations for search results (Yu et al., 2022; Polley et al., 2021; Baier and Kaisers, 2020; Ai et al., 2019; Ai and Narayanan. R, 2021) and showing the impact of explanations on efficiency and trust (Ramos and Eickhoff, 2020; White et al., 2009; Eickhoff et al., 2014), as well as explaining the relevance of search results to users (Coyle and Smyth, 2005; Hearst, 1995). However, there is still a limited understanding of if and when users find search explanations beneficial. Our study explores the potential advantages of search explanations, identify the search objectives in which they prove valuable and ascertain the specific characteristics desired in valuable explanations.
@@ 3.1.2. Targeted user group
No. Search objective Example search queries Perceived helpfulness of explanations
@@
[Search objectives]Table displays the search objectives along with two sample search queries: one provided by us as an example and the other filled in by participants during the survey. Additionally, the table presents the count of participants who found explanations helpful for each search objective.
@@ 3.2.2. Study survey
[Participant demographics]Demographics information of our interview participants.
@@
[…]
[Sample search explanation]Figure showing search explanations provided as an example during the interview. This example illustrates an overall explanation describing how the search engine works and, in addition, explains how relevance influences the ranking of search results.
Figure 1. (a) Figure showing highlighted keywords in the snippet of text accompanying the search result, offering insights into the alignment of the search result with the user’s search query, (b) Figure showing search explanations provided as an example during the interview. This example illustrates an overall explanation describing how the search engine works and, in addition, explains how relevance influences the ranking of search results.
@@ 3.3.2. Interview protocol
Participants were first asked general warm-up questions, including the frequency of their search engine usage and whether they had encountered search explanations on the search results page. They were then prompted to recall situations where they questioned the search results returned by the search engines. Next, we conducted a contextual inquiry task, wherein participants were instructed to search for a specific query on their preferred search engine and share their screen. They were asked to describe their process for selecting search results and provide suggestions for additional information that could enhance their search experience. For this task, we selected two search queries voluntarily provided by participants for various search scenarios in the study survey. These included queries about both non-controversial topics such as ‘best ways to learn investment,’, ‘date ideas in Philly’, etc. and more controversial subjects (if such examples were provided by the participants) such as ‘covid-19 vaccines,’ ‘climate change’, etc. Recall that participants had initially selected these search queries from their browser histories during the study survey, mirroring real-world scenarios where they actively formulate search inquiries based on their information needs. Our goal was to ensure that the chosen queries resonated with participants’ familiarity and interests. During the interview process, some participants chose to search for additional or different search queries in addition to the ones we selected for them. We allowed this flexibility to ensure participants’ comfort and autonomy. Note that the objective of this task is not to determine specific search features needed in particular scenarios but rather to capture the range of potential features that could enhance participants’ overall search experience.
@@ 5.1. Utility of explanations in various search objectives
Table 3. Table providing a summary of the perceived benefits of search result explanations as expressed by the participants along with an example quote.
@@
[Perceived benefits of search explanations]Table providing a summary of the perceived benefits of search result explanations as expressed by the participants along with an example quote.
Table 3. Table providing a summary of the perceived benefits of search result explanations as expressed by the participants along with an example quote.
@@ 5.2.2. Explanations could serve as decision-making tools by assisting in the selection of search results:
Table 4. Table providing a summary of the characteristics desired by the participants in search explanations.
@@
[Desired Characteristics of search explanations]Table providing a summary of the characteristics of search explanations that are desired by the participants.
Table 4. Table providing a summary of the characteristics desired by the participants in search explanations.
@@ 5.2.4. Explanations help in understanding the perspectives or categories shown in the search results:
[Additional details provided by Google and Bing]Table illustrating the comparison of key features shown by Google and Bing alongside each search result.
@@ 6.0.1. Option to access additional information is valued:
All participants acknowledged the value of having the option to access additional information about the search result in both Bing and Google search engines. As P7 noted, “That gives a person the searcher the option of taking the extra time to learn the extra information… and that’s always a good thing to have”. Participants also appreciated the concise presentation of the information presented by search engines, as one participant mentioned, “I like that it’s concise and doesn’t show you everything you want to see..you have to click somewhere else, which I do like because it’s like it makes it less cluttered I guess” (P6).
6.0.2. Broad and obvious explanations are not deemed useful:
@@
Many participants did not find the content of the search explanations, currently only provided by Google, particularly useful, deeming it to be either obvious or too broad. P5 provided an example stating that their getting a specific result due to their location in the USA is not very helpful—“I don’t think the result in the region would be particularly helpful, especially if it’s as broad as the United States”. P2 concurred, expressing, “I feel like a lot of this information doesn’t necessarily help me out with some of those qualities that we were talking about..like the trust, the credibility, the efficiency”.
6.0.3. Contestability is desired in search:
@@
The majority of participants appreciated the functionality provided solely by Google to remove search results and offer feedback.. They saw this feature as a means to contest and influence the search results, granting them a voice in determining the relevance and accuracy of the displayed information. Participants appreciated the agency and empowerment that came with being able to shape the search results according to their preferences and needs. One participant emphasized the significance of user input, expressing the necessity of having such a feature—“I feel like it’s something that’s necessary to have something like this because if they did not include like a send feedback option, it would feel sort of wrong in a way, it would feel more so that like Google is saying, we know what’s right, we know what’s best, and you know, you don’t have a say in that” (P6).
6.0.4. Privacy and personalization settings:
@@
Participants found the privacy and personalization settings provided only by Google to be helpful. The majority of them were not aware that they could change their privacy settings to alter the way Google collects their personal information to personalize the search results for them. Participants suggested that instead of having the privacy settings within the search explanations pop up, “it should be easily accessible and visualizable on the search result page itself” (P3).
6.0.5. Website previews increase search efficiency: